### v4

# Continual Adversarial Defense

Abstract
In response to the rapidly evolving nature of adversarial attacks against visual classifiers on a monthly basis, numerous defenses have been proposed to generalize against as many known attacks as possible. However, designing a defense method that generalizes to all types of attacks is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks that emerge as time goes on. A well-matched approach to the dynamic environment lies in a defense system that continuously collects adversarial data online to quickly improve itself. Therefore, we put forward a practical defense deployment against a challenging threat model and propose, for the first time, the Continual Adversarial Defense (CAD) framework that adapts to attack sequences under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial data. We explore and integrate cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles. Extensive experiments validate the effectiveness of our approach against multiple stages of modern adversarial attacks and demonstrate significant improvements over numerous baseline methods. In particular, CAD is capable of quickly adapting with minimal budget and a low cost of defense failure while maintaining good performance against previous attacks. Our research sheds light on a brand-new paradigm for continual defense adaptation against dynamic and evolving attacks.  [^1] Code at https://github.com/cc13qq/CAD
Figure 1: In a dynamic environment, the target model $f_{t}$ faces a diverse range of attacks at different stages and is protected by the defender. During the $i$-th stage, the defense model leverages few-shot defense budget provided by the web crawler, security department, and users of the target model to adapt to new attacks. The defense model also needs to efficiently use memory to store the cache while retaining knowledge of past attacks.
Introduction
Adversarial attack (Madry et al. 2017) aims to deceive deep neural networks (DNNs) by adding subtle perturbations to input images, seriously jeopardizing the reliability of DNNs, particularly in domains sensitive to security and trust. In order to promote the robustness of DNN, adversarial training (Zhang et al. 2019) and purification techniques (Yoon, Hwang, and Lee 2021) are proposed to defend against adversarial attacks by one-shot training in which the model enters a static phase after a single defense training stage (Zhou and Hua 2024), but often exhibit reduced classification accuracy in adversarial examples and can sacrifice classification capacity in clean images.
Contrary to the one-shot defense assumption, the defense system deployed in real scenarios encounters a dynamic environment characterized by numerous unique attacks with higher aggressivity (Li et al. 2020). A well-matched approach to the dynamic environment lies in a constantly evolving defense strategy that continuously adapts to newly emerging adversarial algorithms. A concurrent continual defense work (Zhou and Hua 2024) treats attacks as a task sequence and tries to train a robust model through lifelong learning. However, poor performance in both adversarial and clean data hinders its deployment in reality and calls into question the reasonableness of the continual defense setting.
To eliminate the insufficiency of the above defense setting, we put forward a practical defense deployment and a challenging threat model. As depicted in Figure 1, the defense system is deployed in the cloud and continuously adapts to new emerging attacks from stage to stage using defense budget (i.e. adversarial examples and their ground-truths) sourced from the web crawler, security department, and users of the target model in real-world scenarios. In reality, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget. When adapted to new emerging attacks, the defense system is permitted to use defense budget of previous stages cached in memory. The defense system also has access to an initial attack and shares the same training data and architecture as the target model. Meanwhile, attacks are initiated under the gray-box setting, wherein they possess knowledge of the classifier’s architecture and have access to training data, but remain unaware of the defense mechanism implemented.
Taking real-world considerations into account, we propose four principles for continual defense: (1) Continual adaptation to new attacks without catastrophic forgetting. As a defender, it is crucial to adeptly adapt to a multitude of new attacks occurring at different stages while preserving knowledge gained from previous encounters. (2) Few-shot adaptation. The escalation in the defense budget signifies a greater potential for attacks on the target model. Therefore, we cannot tolerate an excessive number of attacks before adaptation. (3) Memory-efficient adaptation. Over time, a continuous influx of attacks leads to the accumulation of defense budget, potentially posing memory constraints. In practical terms, we may not have sufficient memory capacity to withstand this pressure. (4) High accuracy in classifying clean and adversarial data. A robust defense should not compromise the interests of those it protects. Hence, ensuring high classification accuracy on both clean and adversarial data is paramount.
In this paper, we propose the first Continual Adversarial Defense (CAD) framework, which aims to defend against evolving attacks in a stage-by-stage manner using few-shot defense budget and efficient memory. In the first stage, we use adversarial data from the initial attack to train an initial defense model which is specifically designed to classify adversarial data, complementing the target model. Drawing inspiration from continual learning (CL), we expand the classification layer to adapt to the new attack as incremental classes and then fine-tune this expanded layer using few-shot defense budget. To address the over-fitting issue that arises from the few-shot fine-tuning, we compress the embedding of previous attacks by generating and assigning virtual prototypes to reserve embedding space in the defense model for future attacks. To optimize memory usage in the data domain, we employ prototype augmentation which allows us to maintain the decision boundary of previous stages without the need to store any budget explicitly. After adaptation of one stage, we refine the classifier to prevent undesirable expansion. Simultaneously, a small model is utilized to ensemble the defense model with the target model by estimating reliable logits for input data, ensuring maximum classification accuracy for both clean and adversarial data.
Based on comprehensive experiments conducted on CIFAR-10 and ImageNet-100, the CAD framework demonstrates strong performance in defending against multi-stage attacks relying on few-shot defense budget while maintaining high accuracy on clean data. To the best of our knowledge, this is the first time to discuss the concept and deployment scenario of continual adversarial defense.
Our main contributions can be summarized as follows:
$\bullet$ To withstand the rapidly evolving adversarial attacks targeting DNNs, we put forward the deployment of continual defense strategy and propose a practical threat model in which diverse attacks emerge at different stages. The defense system must adapt well to these attacks using a few-shot defense budget and efficient memory management, while also retaining knowledge of past attacks and maintaining high performance on both clean and adversarial data.
$\bullet$ We propose, for the first time, the Continual Adversarial Defense (CAD) framework that defends against attacks in the dynamic scenario under four practical principles: continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial data.We leverage cutting-edge techniques including continual learning, few-shot learning, non-exemplar class incremental learning, and ensemble learning to satisfy the principles.
$\bullet$ Extensive experiments validate the effectiveness of CAD against multiple stages of modern adversarial attacks using few-shot budget and efficient memory and demonstrate significant improvements over baseline methods.
Related Work
Adversarial Defense
Researchers have proposed various general defense methods against adversarial attacks (Madry et al. 2017; Wu et al. 2021; Lin et al. 2019; Wang and He 2021). As a standard defense method, adversarial training (Zhang et al. 2020, 2019; Grathwohl et al. 2020; Wang et al. 2023b) aims to enhance the robustness of the target model by training it with adversarial examples. Another branch of adversarial defense involves purifying the data stream (Salman et al. 2020; Hill, Mitchell, and Zhu 2020; Yoon, Hwang, and Lee 2021; Nie et al. 2022) to remove potential adversarial perturbations or noise that could deceive the model. In fact, all of the above methods have shown limitations in robustness, resulting in a lower classification performance on both adversarial and clean data. In this paper, we present the first Continual Adversarial Defense (CAD) framework as a solution to effectively counter a wide spectrum of attacks emerging stage by stage, while ensuring robust performance in classifying clean data.
Continual Learning
Continual learning (CL) (Liu et al. 2022; Kirkpatrick et al. 2017; Wang et al. 2022; Zhou et al. 2022) aims to learn from a sequence of new classes without forgetting old ones and attracts a lot of attention to various computer vision tasks. In recent years, some methods have been aimed at solving the CL problem without relying on preserving data (called no exemplar) (Zhu et al. 2022), and some methods have attempted to learn new classes in a few-shot scenario (Zhang et al. 2021) in which only a small number of new class data is gainable. Attempting to break the one-shot defense assumption, a concurrent work (Zhou and Hua 2024) is proposed to alleviate catastrophic forgetting in a simple continual defense scenario. In this paper, we convert the proposed defense scenario into a few-shot and non-exemplar CL setting, which necessitates the defense mechanism to utilize few-shot defense budget and efficient memory for adaptation.
Threat Model and Defense Principles
The defense system is deployed online and protects the target model against a dynamic environment marked by a plethora of unique attacks. As time progresses, new attacks accompanied by poisoned examples will emerge on the Internet, aiming to exploit vulnerabilities in the target model and achieve breakthroughs. Due to the reality, we propose a practical threat model in which distinct attacks emerge across various stages, requiring the defense system to adapt to evolving threats through the utilization of the defense budget (i.e. adversarial examples and their ground-truths). Typically, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget. The defense budget is crucial as it informs and influences the defender’s response to emerging threats, often discovered during screening or cleanup efforts in the real world, and is provided by web crawlers, security services, and users of the target model. The continual defense scenario is formulated as follows.
Figure 2: Illustration of CAD. Regarding the classes with new attacks as incremental classes, we use continual adaptation to protect the target model while preventing catastrophic forgetting. In the 00-th stage, we reserve the embedding space for future attacks during training of the initial defense model $f_{0}$ to alleviate the overfitting brought by future fine-tuning. Then, a weight estimator model $f_{p}$ is trained for model ensemble. In the $i$-th stage ($i>0$), we freeze the feature extractor $\varphi$ of defense model $f_{i}$ and fine-tune the expanded classifier $g_{c}^{i}$ using few-shot defense budget. To efficiently use memory, we store class-wise prototypes in the prototype cache at each stage and use them in the fine-tuning process. After all, the classifier is refined, and the target model $f_{t}$ and the scalable defense model $f_{i}$ are ensembled by the weight estimator model $f_{p}$ to maintain high performance in classifying clean and adversarial data.
Our scenario is established on the $N$-way $K$-shot classification paradigm. Attacks are launched under the gray-box setting (Taran et al. 2019) in which the attacks know the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space. The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at the 0-th stage. At the $i$-th stage, where $i=1,2,\cdots,T$, new attacks $A_{i}(\cdot)$. are detected and the defender receives a set of budget $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack. Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
Taking reality into account, the defense mechanism should satisfy the following principles:
Principle 1
Continual adaptation to new attacks without catastrophic forgetting.
In the dynamic environment, various new attacks will emerge on the internet as time goes on. Therefore, the defense mechanism must be capable of adapting to a range of new attacks across different time stages while also retaining knowledge of previous ones.
Principle 2
Few-shot adaptation.
The increase in defense budget directly correlates with the frequency of successful attacks on the target model. The defense system must proactively adapt the defense based on few-shot budget to prevent potential security disasters that could arise from delayed action.
Principle 3
Memory-efficient adaptation.
Over time, the ongoing influx of attacks results in accumulating defense budgets, potentially leading to memory constraints. In practical terms, the defender may not have sufficient memory space to accommodate this accumulation. Consequently, storing all received defense budget for adaptation is impractical and should be avoided.
Principle 4
High accuracy in classifying both clean and adversarial data.
The previous defense strategies have led to a sacrifice in the classification accuracy of clean data, resulting in performance degradation. This decline could potentially harm crucial real-world business operations. As a result, a defender must prioritize maintaining performance on both clean and adversarial data.
Continual Adversarial Defense Framework
As shown in Figure 2, Continual Adversarial Defense (CAD) framework consists of continual adaptation to defend the target model while preventing catastrophic forgetting, embedding reservation to alleviate overfitting brought by the few-shot budget, prototype augmentation for memory efficiency, and model ensemble to ensure the classification performance on both clean and adversarial data.
Continual Adaptation to New Attacks
In response to Principle 1, we use continual adaptation to defend against new attacks while preventing catastrophic forgetting. In the 00-th stage, an initial defense model $f_{0}:\mathcal{X}\to\mathbb{R}^{N}$ which consists of a feature extractor $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and a classifier $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ is optimized under full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. The defense model is designed to tackle adversarial examples, complementing the target model. In the $i$-th stage, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to the new attack using the defense budget $\mathcal{A}_{\mathrm{train}}^{i}$.
Inspired by continual learning (CL) (Rebuffi et al. 2017), we regard the classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion. Due to the internal similarity of some adversarial attacks (Wang et al. 2023a), trained with one attack, the defense model can, to some extent, generalize to other attacks, resulting in some unnecessary class expansion. Old attacks occurring in new stages also lead to unnecessary adaptation. Therefore, we conduct model tests and filter out the useless data from the defense budget that can be correctly classified.
Following (Liu et al. 2022), we expand the classifier from $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ to $g_{c}^{1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{2}}$ in the first stage, and expand the classifier $g_{c}^{i}$ from $\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$ in the $i$-th stage. The parameter of the expanded classifier is composed of the parameter of the old classifier and the newly initialized parameter: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$. Meanwhile, the ground truth of $\mathcal{A}_{\mathrm{train}}^{i}$ is rewritten corresponding to the incremental classes with the new attack  [^2] We omit the networks’ parameter in formulas:
$y^{i}=y+N\times{i}.$ (1)
The initial defense model $f_{0}$ will be trained in the Embedding Reservation Section. After this, we freeze the feature extractor $\varphi$ and fine-tune the expanded classifier part for the new attack $A_{i}$ using the few-shot budget $\mathcal{A}_{\mathrm{train}}^{i}$ in the Prototype Augmentation Section.
For evaluation, we are just focusing on which class the image is placed in rather than being threatened by which attack. Thus the prediction for an instance $\mathbf{x}$ is:
$y_{\mathrm{pred}}=(\arg\max f_{i}(\mathbf{x}))\%N$ (2)
Embedding Reservation for Few-Shot Adaptation
A problem brought by the few-shot budget in Principle 2 is over-fitting which is common in few-shot learning (Snell, Swersky, and Zemel 2017). To tackle this issue, we choose to reserve embedding space for future attacks by generating and assigning virtual prototypes, squeezing the embedding space of previous attacks.
First, before training the defense model $f_{0}$ at the 00-th stage, several virtual prototypes $P_{v}=[\mathbf{p}_{v}^{1},\cdots,\mathbf{p}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier and treated as ”virtual classes” (Zhang et al. 2021), where $V=N\times T$ is the number of virtual classes, i.e. the reserved classes for future attacks. Therefore the output of current model is $f_{0}(\mathbf{x})=[W^{0},P_{v}]^{\top}\varphi(\mathbf{x})$.
Second, virtual instances are constructed by manifold mixup (Verma et al. 2019):
$\mathbf{z}=h_{2}(\lambda h_{1}(\mathbf{x}_{\mathrm{adv},r}^{0})+(1-\lambda)h_{1}(\mathbf{x}_{\mathrm{adv},s}^{0})),$ (3)
where $\mathbf{x}_{\mathrm{adv},r}$ and $\mathbf{x}_{\mathrm{adv},s}$ belong to different classes $r$ and $s$, and $\lambda\sim\mathrm{B}(\alpha,\beta)$ is a trade-off parameter the same as (Zhou et al. 2022). $h_{1}$ and $h_{2}$ are decoupled hidden layers of feature extractor i.e. $\varphi(\mathbf{x})=h_{2}\circ h_{1}(\mathbf{x})$.
Finally, the embedding space reserving is conducted by training $f_{0}$ with the following loss:
$\mathcal{L}_{0}=$ $\gamma F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0}),\hat{y})+F_{\mathrm{ce}}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0})+$
$+$ $\gamma F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{z}),\hat{y}),\hat{\hat{y}})+F_{\mathrm{ce}}(f_{0}(\mathbf{z}),\hat{y}).$ (4)
where $\hat{y}=\mathrm{argmax}_{j}\mathbf{p}_{v}^{j\top}\varphi(\mathbf{x}_{\mathrm{adv}}^{0})+N$ is the virtual class with maximum logit, acting as the pseudo label. $\hat{\hat{y}}=\mathrm{argmax}_{k}\mathbf{w}_{k}^{0\top}\mathbf{z}$ is the pseudo label among current known classes. $\gamma$ is a trade-off parameter, $F_{\mathrm{ce}}$ represents the standard cross-entropy loss (Zhang and Sabuncu 2018), and function $\mathsf{Mask}(\cdot)$ masks out the logit corresponding to the ground-truth:
$\mathsf{Mask}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0})=f_{0}(\mathbf{x}_{\mathrm{adv}}^{0})\otimes(\mathbf{1}-\mathrm{onehot}(y^{0}))$ (5)
where $\otimes$ is Hadamard product and $\mathbf{1}$ is an all-ones vector.
In Eq. Embedding Reservation for Few-Shot Adaptation, it pushes an instance away from the reserved virtual classes (item 1) while towards known classes (item 2), and forces a virtual instance away from known class (item 3) while towards virtual class (item 4). Trained with $\mathcal{L}_{0}$, the embedding of initial benign classes will be more compact, and the embedding spaces for virtual classes will be reserved (Zhou et al. 2022). The reserved space allows the defense model to be adapted more easily in the future and alleviates overfitting brought by few-shot budget.
Prototype Augmentation for Memory Efficiency
In response to Principle 3, we use prototype augmentation to efficiently manage memory.
When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased (Zhu et al. 2021). Many CL methods store a fraction of old data to jointly train the model with current data to address this issue (Rebuffi et al. 2017). However, preserving old budget may also lead to memory shortage. For memory efficiency, we adopt prototype augmentation (Zhu et al. 2022) to maintain the decision boundary of previous stages, without saving any budget. We memorize one prototype in the deep feature space for each class with each attack, and over-sample (i.e., $U_{B}$) prototypes $\mathbf{p}_{B}$ and ground-truths $y_{B}$ to the batch size, achieving the calibration of the classifier:
$\mathbf{p}_{B}=U_{B}(P_{e}),\mathcal{L}_{\mathrm{proto}}=F_{\mathrm{ce}}(\mathbf{p}_{B},y_{B})$ (6)
where $P_{e}=\{\mathbf{p}_{e}^{j}\}_{j=0}^{N\times i}$ is the set of class-wise average embedding (i.e., prototype cache) generated by feature extractor $f_{e}^{i}$ following ProtoNet (Snell, Swersky, and Zemel 2017):
$\mathbf{p}_{e}^{j}=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}(y_{k}^{i}=j)\varphi(\mathbf{x}_{\mathrm{adv},k}^{i})$ (7)
where $\mathbb{I}(\cdot)$ is the indicator function, and $K$ is the amount of budget for each class.
At each stage, the defense model is adapted using the following loss:
$\mathcal{L}_{\mathrm{ce}}=F_{\mathrm{ce}}(f_{i}(\mathbf{x}_{\mathrm{adv}}^{i}),y^{i}),\quad\mathcal{L}_{f}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{proto}}$ (8)
Model Ensemble for Clean Data Classification
To maintain high classification accuracy on both clean and adversarial data mentioned in Principle 4, we propose the model ensemble as the last part of CAD. Ensemble adversarial training (EAT) (Tramèr et al. 2018) which trains a robust model using adversarial samples generated by the target model is a simple yet effective way to defend against adversarial attack under the gray-box setting while maintaining the classification performance of clean data. We extend EAT to our scenario by training a small weight estimator model $f_{p}$ to fuse the logits of models.
We adopt self-perturbation (Wang et al. 2023a) to train the weight estimator at the first stage. Agnostic to any of the attacks, a self-perturbation-trained model is able to distinguish images from benign and adversarial. The weight estimator $f_{p}$ outputs a weight vector $\mathbf{w}\in\mathbb{R}^{2}$ to ensemble the defense model and the target model:
$\mathrm{logit}_{i}(\mathbf{x})=\mathbf{w}\cdot[f_{t}(\mathbf{x}),f_{i}(\mathbf{x})]^{\top}$ (9)
In this way, both clean and adversarial images are correctly classified.
For the overall algorithm of the CAD framework, details about self-perturbation, and limitations of CAD, please refer to the supplementary materials.
Experiment
To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies comparing CAD with baselines from three research streams. For evaluation and analysis, we use two metrics in this section: (1) Classification accuracy (Acc) against each attack after adaptation for defense performance evaluation. (2) Average accuracy of all attacks occurring within and before each stage, i.e. average incremental accuracy (AIAcc) (Rebuffi et al. 2017), for assessing the knowledge retention of the model.
Datasets
Experiments are conducted on CIFAR-10 and ImageNet-100. CIFAR-10 (Krizhevsky, Hinton et al. 2009) contains 5000 images for training and 1000 images for testing in each of the 10 classes. ImageNet-100 (Tian, Krishnan, and Isola 2020) is a subset of ImageNet (Deng et al. 2009) and contains 100 classes with 1000 images for training and 100 images for testing in each class. Images from CIFAR-10 and ImageNet-100 are resized to $32\times 32$ and $224\times 224$ respectively. Before training, the above datasets endured data augmentation including horizontal flip and random crop.
Attack Algorithms
We choose PGD-$l_{\infty}$ (Madry et al. 2017) as the initial attack. After that, we select 8 adversarial attack methods under $l_{\infty}$: BIM (Kurakin, Goodfellow, and Bengio 2016), RFGSM (Tramèr et al. 2018), MIM (Dong et al. 2018), DIM (Wu et al. 2021), NIM (Lin et al. 2019), SNIM (Lin et al. 2019), VNIM and VMIM (Wang and He 2021) to compose the attack pool in defense performance evaluation. The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following (Zhang et al. 2019). As for evaluating the capability of CAD in migrating catastrophic forgetting, $l_{\infty}$ attack VMIM and SNIM, $l_{2}$ attack CW (Carlini and Wagner 2017) and DeepFool (Moosavi-Dezfooli, Fawzi, and Frossard 2016), $l_{1}$ attack EAD (Chen et al. 2018) and EADEN (Chen et al. 2018), and $l_{0}$ attack OnePixel (Su, Vargas, and Sakurai 2019) and SparseFool (Modas, Moosavi-Dezfooli, and Frossard 2019) compose the attack pool. Details of attack parameters are given in the supplementary.
Stage:Attack 0:PGD 1:SNIM 2:BIM 3:RFGSM 4:MIM 5:DIM 6:NIM 7:VNIM 8:VMIM Clean
None-defense 0.2 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.0 89.8
TRADES (Zhang et al. 2019) 46.5 45.5 49.6 46.2 48.0 47.7 46.7 48.2 49.5 76.2
GAIRAT (Zhang et al. 2020) 59.8 60.2 58.4 59.2 58.8 58.9 58.9 59.7 59.8 74.1
RPF (Dong and Xu 2023) 56.4 49.9 51.3 56.6 54.4 52.6 53.9 53.3 50.1 69.4
DMAT (Wang et al. 2023b) 64.3 65.1 66.2 65.9 65.4 64.4 64.1 65.8 65.6 75.9
ADP (Yoon, Hwang, and Lee 2021) 77.2 76.9 71.8 71.8 75.6 75.9 76.2 76.0 72.3 74.1
DiffPure (Nie et al. 2022) 53.2 53.0 52.4 55.8 53.9 53.4 52.6 54.7 52.3 73.6
CAD(ours) 83.7 83.4 82.0 82.5 83.5 82.0 81.9 82.2 81.9 88.1
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 83.5 83.9 84.0 82.6 84.3 83.8 84.2 82.2 81.4 87.8
Table 1: Classification accuracy(%) against attacks($l_{\infty}$) on ImageNet-100. Volumn ”Clean” represents the standard accuracy of clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD in case no constraints on memory and defense budget, and is considered to be the empirical upper bound marked with a gray background.
Stage:Attack 0:PGD 2:SNIM 3:VNIM 4:CW 5:DeepF. 6:EAD 7:EADEN 8:OnePixel 9:SparseF.
None-defense 0.6 1.3 0.0 0.0 8.6 0.0 0.2 39.1 29.7
EWC (Kirkpatrick et al. 2017) 95.5 68.2 67.0 23.7 23.6 23.7 26.9 25.9 22.2
SSRE (Zhu et al. 2022) 89.1 82.5 78.1 65.8 58.4 50.6 47.5 42.4 37.8
BEEF (Wang et al. 2022) 94.9 83.1 84.7 82.3 77.9 74.2 73.8 72.6 64.5
FACT (Zhou et al. 2022) 94.5 87.2 86.4 81.5 80.9 78.6 78.4 78.2 74.6
CAD(ours) 94.5 89.1 90.5 84.1 84.8 81.1 82.1 82.8 78.2
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.1 95.2 94.9 95.3 94.6 96.0 95.2
Table 2: Continual defense performance (average incremental accuracy(%)) against attacks under different norms $l_{p=0,1,2,\infty}$ on CIFAR-10. DeepF. and SparseF. denote DeepFool and SparseFool respectively.
Implementation Details
Both the target model and defense model use ResNet (He et al. 2016) as the backbone network in our method. Specifically, we use WideResNet-28-10 for CIFAR-10 following (Grathwohl et al. 2020) and ResNet-50 for ImageNet-100 following (Laidlaw, Singla, and Feizi 2020). A 4-layer ConvNet is used as the weight estimator model. The cosine classifier (Gidaris and Komodakis 2018) is adopted as the classifier of the defense model. We utilize Torchattacks (Kim 2020) to generate adversarial images. Shots of defense budget per class is set to $K=10$. The number of stages is set to $T=8$ since there are 9 attacks (PGD is the initial attack). Parameters of Beta distribution are set to $\alpha=\beta=2$ and the trade-off parameter for $\mathcal{L}_{0}$ is set to $\gamma=0.01$ following (Zhou et al. 2022). Epochs for training and fine-tuning are set to $100$ and $4$ respectively.
Defense Baselines
We first compare with adversarial training methods TRADES (Zhang et al. 2019), GAIRAT (Zhang et al. 2020), RPF (Dong and Xu 2023) and DMAT (Wang et al. 2023b), and purification methods ADP (Yoon, Hwang, and Lee 2021) and DiffPure (Nie et al. 2022) in defense performance evaluation. As for evaluating the capability of CAD in migrating catastrophic forgetting, we compare to CL methods EWC (Kirkpatrick et al. 2017), SSRE (Zhu et al. 2022), BEEF (Wang et al. 2022), and FACT (Zhou et al. 2022). Besides, we propose another variant of our method as the empirical upper bound, denoted as premium-CAD. For premium-CAD, there is no constraint of memory and no concerns about the amount of defense budget. We train additional attack-specific defense models using abundant defense budget in each stage, and ensemble them with the target model by the weight estimator. For details about the implementation of defense baseline and premium-CAD, please refer to supplementary materials.
Comparisons to Baselines
To evaluate the defense performance, we first compared CAD to defense baseline methods in defending against 9 adversarial attacks. For CAD, the defense environment is dynamic, with attacks occurring at different stages. The other baseline methods operate in a static defense environment that does not differentiate the order of attacks. As depicted in Table 1, despite the small scale of the defense budget, CAD demonstrates strong adaptability to each type of attack. When compared with the top baseline DMAT and ADP on ImageNet-100, CAD outperforms them by approximately $6\%$ for each attack. This underscores its capacity for continuous adaptation in response to emerging attacks. Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $88.1\%$ accuracy, while the target model stands at $89.8\%$. This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images.
(a) PGD
(b) CW
(c) EAD
(d) OnePixel
Figure 3: T-SNE diagrams of features of test data encoded by the defense model after all stages. Each color represents a class of CIFAR10. After several steps of adaptation, CAD retains the capability to categorize the attacked data.
To assess the effectiveness of CAD in retaining knowledge, we compare it to CL baseline methods on CIFAR-10 in a dynamic scenario formed by attacks under different norms $l_{p=0,1,2,\infty}$. As depicted in Table 2, CAD maintains robust defense performance despite changes in attack types. However, as the variety of attack types increases, the defense performance gradually diminishes. This indicates that defending against an expanding array of attack types presents a more challenging task for continual defense. Compared with the top baseline method FACT, CAD consistently outperforms it by an average of approximately $3.4\%$, highlighting CAD’s effectiveness in mitigating catastrophic forgetting.
It is noteworthy that premium-CAD surpasses all baseline methods. Given sufficient defense budget and memory resources, premium-CAD emerges as the superior choice.
Ablation Study
As argued, embedding reservation facilitates easier adaptation of the defense model and mitigates overfitting. Prototype augmentation preserves the decision boundary of previous stages, while model ensemble ensures robust classification performance for both clean and adversarial images.
Stage:Attack 0:PGD 1:CW 2:EAD 3:OnePixel Clean
w/o $\mathcal{L}_{0}$ 94.5 82.4 76.1 71.8 95.8
w/o $\mathcal{L}_{\mathrm{proto}}$ 94.4 89.5 80.0 78.5 95.8
w/o $f_{p}$ 94.5 89.9 86.6 85.3 55.9
CAD(ours) 94.5 89.9 86.6 85.3 95.8
Table 3: AIAcc(%) on CIFAR-10 for ablation.
Therefore, we proceed to conduct an ablation study using 4 attacks PGD($l_{\infty}$), CW($l_{2}$), EAD($l_{1}$), and OnePixel($l_{0}$) to validate the efficacy of each component. As shown in Table 3, CAD without $\mathcal{L}_{0}$ and $\mathcal{L}_{\mathrm{proto}}$ experiences a decrease of $13.5\%$ and $6.8\%$ respectivily in the last stage. The absence of model ensemble leads to the inability to uphold performance on clean images.
Additionally, we provide T-SNE diagrams of features extracted from adversarial test data in Figure 3 to visualize the defense capability of CAD after adaptation. These features are encoded by the defense model after the last stage of adapting to OnePixel attack. From (a) to (d) in Figure 3, the clustering of features becomes less and less discernible, indicating that the difficulty of few-shot adaptation increases with the introduction of new attack types. After adaptation, CAD maintains its defense performance against previous attacks, illustrating its ability to mitigate catastrophic forgetting as demanded in Principle 1.
Multiple Attacks in One Stage
Stage 0 1 2 3
$\mathrm{number}=1$ 94.5 95.4 93.4 94.7
$\mathrm{number}=2$ 94.5 94.4 93.5 92.8
$\mathrm{number}=3$ 94.5 93.9 93.6 92.5
Table 4: Acc(%) against multiple attacks in one stage.
In practice, it is common for multiple attacks to occur simultaneously within a single stage. To assess CAD’s adaptability in such scenarios, on CIFAR-10, we evaluate the classification accuracy when faced with 1, 2, and 3 concurrent attacks per stage. As presented in Table 4, when the number of attacks increases to $2$ and $3$, the classification accuracy against these attacks still remains at a high level: $92.8\%$ and $92.5\%$ at stage $3$ respectively. The results indicate that CAD continues to perform effectively even when confronted with multiple simultaneous attacks.
Number of Shots for Adaptation
Stage:Attack 0:PGD 1:CW 2:EAD 3:OnePixel
$K=15$ 94.5 90.0 88.4 85.5
$K=10$ 94.5 89.9 86.6 85.3
$K=5$ 94.5 89.4 86.1 84.6
$K=1$ 94.5 87.9 84.6 83.4
Table 5: AIAcc(%) with different budget scale.
As indicated in Principle 2, the defense adaptation relies on few-shot defense budget. To investigate the influence of the budget shots $K$, we conducted tests with $K=15,10,5,1$ on CIFAR-10. The results presented in Table 5 illustrate that defense performance diminishes as $K$ decreases. Nevertheless, even with reduced $K$, CAD continues to efficiently defend against adversarial attacks ($83.4\%$ when $K=1$ at stage $3$), highlighting its strong capability in few-shot scenarios.
Memory Allowance for Adaptation
As outlined in Principle 3, we need to employ memory efficiently for continual defense. In order to demonstrate CAD’s effectiveness in conserving memory, we conducted tests to measure the occupied memory space and model size at each
Cache Size Model Size
Stage 1 2 3 1 2 3
CAD 50.3K 75.5K 100.4K 141M 141M 141M
$\mathrm{\textbf{CAD}}^{\dagger}$ 257M 377M 504M 280M 419M 558M
Table 6: Comparison of the Cache Size and Model Size (Byte) between CAD and premium-CAD.
stage on CIFAR-10. As depicted in Table 6, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size. This underscores our approach’s ability to utilize efficient memory.
Conclusion
The operational environment for defense systems is inherently dynamic, making it unrealistic to expect any defense method to effectively address all types of attacks. In response to this challenge, we introduce the first Continual Adversarial Defense (CAD) framework. CAD is engineered to collect few-shot defense budget from the Internet and dynamically adapt to a diverse range of attacks as they emerge in sequential stages. In consideration of practical applicability, we have formulated four principles that include continual adaptation without catastrophic forgetting, few-shot budget adaptation, memory-efficient adaptation, and high accuracy for both clean and adversarial data, and CAD demonstrates strong performance when operating under these principles. Through extensive experiments, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.
Appendix
Algorithm 1 Continual Adversarial Defense
Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, initial attack $A_{0}$, target model $f_{t}$, and few-shot budget $\mathcal{A}_{\mathrm{train},K}^{i}$ generated by attack $A_{i}$ in stage $i=1,\cdots,T$.
Output: Defense model ensemble.
1:  Generate initial adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}$ using $A_{0}$.
2:  Train weight estimator $f_{p}$ on $\mathcal{D}_{\mathrm{train}}$ using self-perturbation.
3:  Rewrite the ground-truth of $\mathcal{A}_{\mathrm{train}}^{i}$ using Eq.1.
4:  Optimize defense model $f_{0}$ on $\mathcal{A}_{\mathrm{train}}^{0}$ using Eq.4.
5:  Generate prototypes of $A_{0}$ using Eq.7 and add them to prototype set $P_{e}$.
6:  for $i$ in $1,\cdots,T$ do
7:     Fine-tune the part of the classifier preserved for the new attack using Eq.8.
8:     Generate prototypes of $A_{i}$ using Eq.7 and add them to prototype set $P_{e}$.
9:     Refine the classifier using cached prototypes $P_{e}$.
10:     Ensemble the defense model $f_{i}$ and the target model $f_{t}$ using Eq.9.
11:  end for
Appendix A More Details about CAD
The overall algorithm of the CAD framework is presented in Algorithm 1.
Model Ensemble Base on Self-Perturbation
In CAD, we use clean images $\mathcal{D}_{\mathrm{train}}$ and self-perturbation (Wang et al. 2023a) to train a weight estimator to ensemble the target model and the defense model. We craft half of the real images to pseudo adversarial images as negative samples and label other real images as positive samples. Then, we train a 4-layer ConvNet as the weight estimator $f_{p}$ in a binary classification manner. In this section, we introduce the details of self-perturbation and model ensemble.
Algorithm 2 Self-perturbation
Input: A empty perturbation matrix $\boldsymbol{\eta}^{p}\in\mathbb{R}^{H\times W\times 3}$ with the same shape of real image $\mathbf{x}^{r}$.
Parameter: Max perturbation magnitude $\epsilon$, pattern mode.
Output: Self-perturbed image $\mathbf{x}^{p}$.
1:  A random direction matrix $R=\{\vec{r}_{ij}\}\in\mathbb{R}^{H\times W\times 3}$.
2:  for $\vec{\eta}_{ij}$ in $\boldsymbol{\eta}^{p}$ do
3:     Select random noise value $\alpha$.
4:     if pattern mode is ‘point-wise’ then
5:        $\vec{\eta}_{ij}:=\alpha\cdot\vec{r}_{ij}$.
6:     else if pattern mode is ‘block-wise’ then
7:        Select a random neighborhood $A_{ij}$ of $\vec{\eta}_{ij}$.
8:        for $\vec{\eta}_{ijk}$ in $A_{ij}$ do
9:           $\vec{\eta}_{ijk}:=\alpha\cdot\vec{r}_{ij}$.
10:        end for
11:     end if
12:  end for
13:  Clip perturbation $\boldsymbol{\eta}^{p}$ using Equation 10.
14:  Generate self-perturbed image $\mathbf{x}^{p}$ using Equation 11.
15:  return $\mathbf{x}^{p}$
As presented in Algorithm 2, we perturb each point in the point-wise pattern and each block in the block-wise pattern in a stochastic direction, where blocks are random neighborhoods of a set of scattered points. The generated perturbation image $\boldsymbol{\eta}^{p}$ is constrained in $l^{\infty}$ norm, and clipped according to $\epsilon$,
$\boldsymbol{\eta}^{p}=\mathrm{Clip}_{[-\epsilon,\epsilon]}(\boldsymbol{\eta}^{p}).$ (10)
A self-perturbed image is calculated as
$\mathbf{x}^{p}=\mathbf{x}^{r}+\boldsymbol{\eta}^{p}.$ (11)
Trained with self-perturbation, the weight estimator $f_{p}$ is able to distinguish between the clean and adversarial images. The output logit $\mathbf{w}$ of $f_{p}$ represents the probability of an image to be clean or adversarial. Therefore, we use $\mathbf{w}$ to ensemble the target model and the defense model:
$\mathrm{logit}_{i}(\mathbf{x})=\mathbf{w}\cdot[f_{t}(\mathbf{x}),f_{i}(\mathbf{x})]^{\top}$ (12)
In this way, both the clean images and the adversarial images are correctly classified.
Stage 0 1 2 3 4 5 6 7 8 -
Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
None-defense 0.6 1.3 0.0 0.0 0.0 0.0 0.2 0.0 0.0 96.4
TRADES (Zhang et al. 2019) 55.8 60.2 57.4 55.9 55.8 57.1 57.1 58.9 57.3 84.8
JEM (Grathwohl et al. 2020) 55.8 62.0 46.5 52.3 52.4 51.1 51.4 57.0 48.2 91.7
PAT (Laidlaw, Singla, and Feizi 2020) 45.5 45.1 43.8 45.5 45.3 45.2 45.2 43.9 45.4 69.3
GAIRAT (Zhang et al. 2020) 66.9 70.4 66.4 66.9 66.7 66.5 66.5 68.6 66.4 89.4
FastAdv (Jiang et al. 2023) 37.1 39.6 36.8 37.0 37.7 37.7 54.6 36.4 37.6 75.2
RPF (Dong and Xu 2023) 60.9 52.8 52.2 58.9 54.9 57.6 57.5 53.5 52.6 83.5
DMAT (Wang et al. 2023b) 71.7 76.0 73.1 71.7 71.7 73.0 73.0 75.5 73.0 92.4
EBM (Hill, Mitchell, and Zhu 2020) 75.2 72.9 73.0 76.6 76.5 72.4 73.9 74.0 73.8 86.8
ADP (Yoon, Hwang, and Lee 2021) 85.4 80.5 79.3 84.4 85.1 84.1 83.7 81.0 79.1 80.5
DiffPure (Nie et al. 2022) 72.5 70.2 69.8 70.1 72.2 73.4 73.5 72.7 69.4 89.3
CAD(ours) 94.5 95.4 93.4 94.7 95.0 93.7 93.7 95.2 93.7 95.8
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.6 94.1 94.6 94.3 95.1 95.1 95.7 96.0
Table 7: Classification accuracy(%) against various attacks($l_{\infty}$) on CIFAR-10. Volumn ”Clean” represents the standard accuracy of clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD that is not constrained by efficient memory and few-shot defense budget.
Appendix B Premium-CAD
We propose another variant of our method denoted as premium-CAD in the case of no constraint of memory and no concerns about the amount of feedback. In each stage, we train an attack-specific defense model using abundant feedback and ensemble the target model and all the defense models by a stage-evolved weight estimator.
Attack-Specific Defense Model
In stage 00, the initial defense model is optimized under the full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. In stage $i$, another defense model is trained using abundant feedback $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)\}_{N\times K}$ where $N$ is the number of classes in dataset $\mathcal{D}_{\mathrm{train}}$ and $K$ is the amount of feedback per class. The scale of feedback is set to $K=10$ in CAD and $K=1000$ in premium-CAD.
In this way, we have $i+1$ defense models in stage $i$. The defense model is designed to tackle adversarial examples from a specific attack, complementing each other.
Algorithm 3 Premium Continual Adversarial Defense
Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, abundant feedback $\mathcal{A}_{\mathrm{train}}^{i}$ generated by attack $A_{i}$ for stage $i=1,\cdots,T$.
Output: Defense model ensemble.
1:  for $i$ in $1,\cdots,T$ do
2:     Train an attack-specific defense model $f_{i}$.
3:     Train a stage-evolved weight estimator $f_{p}^{i}$.
4:     Ensemble the defense models $[f_{0},f_{1},\cdots,f_{i}]$ and the target model $f_{t}$ using Equation 14.
5:  end for
Stage 0 1 2 3 4 5 6 7 8
Attack VMIM VNIM NIM DIM MIM RFGSM BIM SNIM PGD
CIFAR 94.8 94.6 93.4 93.5 94.2 93.4 92.5 92.2 93.3
ImageNet 82.5 82.1 82.0 82.0 92.4 82.2 81.9 82.3 81.7
Table 8: Classification Acc (%) against changed attack sequence.
Stage-Evolved Weight Estimator
To ensemble the target model $f_{t}$ and defense models $[f_{0},f_{1},\cdots,f_{i}]$, we use a joint dataset $\mathcal{D}_{J}^{i}=\bigcup_{i=0,1,\cdots,i}\mathcal{A}_{\mathrm{train}}^{i}\cup\mathcal{D}_{\mathrm{train}}$ to train the weight estimator $f_{p}^{i}$ in stage $i$. Feed an image to the weight estimator $f_{p}^{i}$, it output a weight vector $\mathbf{w}^{i}\in\mathbb{R}^{i+2}$ to ensemble models:
$\mathbf{w}^{i}=f_{i}(\mathbf{x})$ (13)
The final output for premium-CAD is:
$\mathrm{logit}_{i}(\mathbf{x})=\mathbf{w}^{i}\cdot[f_{t}(\mathbf{x}),f_{0}(\mathbf{x}),f_{1}(\mathbf{x}),\cdots,f_{i}(\mathbf{x})]^{\top}$ (14)
The overall algorithm of premium-CAD framework is presented in Algorithm 3. The loss function for training attack-specific defense models and stage-evolved weight estimators is Cross-Entropy Loss.
Appendix C More Implementation Details
Adversarial Attack
In this work, all adversarial attacks are implemented using TorchAttacks (Kim 2020). The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following (Zhang et al. 2019). Iterative steps and step size of iterative attacks are set to $10$ and $2/255$ respectively. The number of perturbed pixels of OnePixel (Su, Vargas, and Sakurai 2019) is set to $50$. The other parameters are the default ones in TorchAttacks.
Defense Baselines
The adversarial training baselines are implemented using RobustBench (Croce et al. 2021), while purification baselines are implemented using open-source code and model checkpoints. CL baselines are implemented using PyCIL (Zhou et al. 2023). The backbone models of both adversarial training methods and CL methods are standardized as WideResNet-28-10 for CIFAR-10 and ResNet-50 for ImageNet-100. CL models are trained in the same settings as in CAD. Specifically, the training epochs and the amount of feedback per class for training CL models are set to 100 and K=10, respectively.
Appendix D More Experimental Results
Continual Adversarial Defense on CIFAR-10
We compare our CAD with more adversarial defense baselines on CIFAR-10. As presented in Table 7, both CAD and premium-CAD demonstrate strong adaptability to each type of attack. When compared with the top baseline methods DMAT (Wang et al. 2023b) and ADP (Yoon, Hwang, and Lee 2021), CAD outperforms them by approximately $10\%$ for each attack. As for clean images, CAD achieves $95.8\%$ accuracy while the target model stands at $96.4\%$.
Change the sequential of attacks
For different attack sequences, we reverse the sequential of attacks and conduct an experiment. As shown in Table 8, when the attack sequence changes, CAD maintains a consistent defense performance, indicating our CAD is agnostic to the attack order.
Change the intensity of attacks
Stage 0 1 2 3 4 5 6
ImageNet-$\epsilon/255$ $4$ $6$ $8$ $10$ $12$ $14$ $16$
Acc 83.7 83.6 83.3 83.3 83.4 82.2 82.5
CIFAR-$\epsilon/255$ $8$ $10$ $12$ $14$ $16$ $18$ $20$
Acc 94.5 94.4 94.4 93.9 94.0 93.2 93.4
Table 9: Classification Acc (%) against various attack intensity of PGD($l_{\infty}$).
For different attack intensities, we conduct experiments using PGD($l_{\infty}$) with various $\epsilon$. As shown in Table 9, when the attack intensity changes, CAD maintains a consistent defense performance, indicating our CAD is agnostic to the attack intensity.
Averaged accuracy when the number of classes increases
We present curves depicting the relationship between averaged accuracy and the number of classes in Figure 4. As the number of classes increases, CAD continues to perform well on both the new attack and previous attacks.
Figure 4: Average Acc with 10 classes and 100 classes per stage for CIFAR10 and ImageNet100.
Appendix E Limitations and Future Work
The increase in attacks denotes the expansion of the defense model classifier, especially when tackling large datasets such as the original ImageNet. Although CAD refines the defense budget before each stage’s classifier expansion, too many classes can lead to an undesirable decrease in classification performance. In the future, we will focus on this issue and refine the classifier to guarantee the consistent performance of the defense model.