### v5

# Continual Adversarial Defense

Abstract
In response to the rapidly evolving nature of adversarial attacks against visual classifiers on a monthly basis, numerous defenses have been proposed to generalize against as many known attacks as possible. However, designing a defense method that generalizes to all types of attacks is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks that emerge as time goes on. A well-matched approach to the dynamic environment lies in a defense system that continuously collects adversarial data online to quickly improve itself. Therefore, we put forward a practical defense deployment against a challenging threat model and propose, for the first time, the Continual Adversarial Defense (CAD) framework that adapts to attack sequences under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial data. We explore and integrate cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles. Extensive experiments validate the effectiveness of our approach against multiple stages of modern adversarial attacks and demonstrate significant improvements over numerous baseline methods. In particular, CAD is capable of quickly adapting with minimal budget and a low cost of defense failure while maintaining good performance against previous attacks. Our research sheds light on a brand-new paradigm for continual defense adaptation against dynamic and evolving attacks.
Figure 1: In a dynamic setting, the target model $f_{t}$ encounters evolving attacks, safeguarded by the defender. At stage $i$, the defense model adapts to new attacks using few-shot defense budget provided by the web crawler, security department, and users of the target model, while efficiently managing memory to retain past attack knowledge.
1 Introduction
Adversarial attack [23] aims to deceive deep neural networks (DNNs) by adding subtle perturbations to input images, seriously jeopardizing the reliability of DNNs, particularly in domains sensitive to security and trust. In order to promote the robustness of DNN, adversarial training [40] and purification techniques [38] are proposed to defend against adversarial attacks by one-shot training in which the model enters a static phase after a single defense training stage [44], but often exhibit reduced classification accuracy in adversarial examples and can sacrifice classification capacity in clean images.
Unlike the one-shot defense assumption, real-world defense systems face dynamic environments with increasingly aggressive and diverse attacks [19]. A well-matched approach to the dynamic environment lies in a constantly evolving defense strategy that continuously adapts to newly emerging adversarial algorithms [3]. However, current model adaptation defenses [2] focus on continuously changing defense strategies when under attack rather than leveraging past defense experiences, leading to suboptimal performance. A concurrent continual defense work [44] treats attacks as a task sequence and tries to train a robust model through lifelong learning. However, poor performance in both adversarial and clean data hinders its deployment in reality and calls into question the reasonableness of the continual defense setting.
To eliminate the insufficiency of the above defense setting, we put forward a practical defense deployment and a challenging threat model. As depicted in Figure 1, the defense system is deployed in the cloud and continuously adapts to new emerging attacks from stage to stage using defense budget (i.e. adversarial examples and their ground-truths) sourced from the web crawler, security department, and users of the target model in real-world scenarios. In reality, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget. When adapted to new emerging attacks, the defense system is permitted to use defense budget of previous stages cached in memory, to maintain the knowledge of previous attacks. The defense system also has access to an initial attack and shares the same training data and architecture as the target model. Meanwhile, attacks are initiated under the gray-box setting, wherein they possess knowledge of the classifier’s architecture and have access to training data, but remain unaware of the defense mechanism implemented.
Taking real-world considerations into account, we propose four principles for continual defense: (1) Continual adaptation to new attacks without catastrophic forgetting. As a defender, it is crucial to adeptly adapt to a multitude of new attacks occurring at different stages while preserving knowledge gained from previous encounters. (2) Few-shot adaptation. The escalation in the defense budget signifies a greater potential for attacks on the target model. Therefore, we cannot tolerate an excessive number of attacks before adaptation. (3) Memory-efficient adaptation. Over time, a continuous influx of attacks leads to the accumulation of defense budget, potentially posing memory constraints. In practical terms, we may not have sufficient memory capacity to withstand this pressure. (4) High accuracy in classifying clean and adversarial data. A robust defense should not compromise the interests of those it protects. Hence, ensuring high classification accuracy on both clean and adversarial data is paramount.
In this paper, we propose the first Continual Adversarial Defense (CAD) framework, which aims to defend against evolving attacks in a stage-by-stage manner using few-shot defense budget and efficient memory. In the first stage, we use adversarial data from the initial attack to train an initial defense model which is specifically designed to classify adversarial data, complementing the target model. Drawing inspiration from continual learning (CL), we expand the classification layer to adapt to the new attack as incremental classes and then fine-tune this expanded layer using few-shot defense budget. To address the over-fitting issue that arises from the few-shot fine-tuning, we compress the embedding of previous attacks by generating and assigning virtual prototypes to reserve embedding space in the defense model for future attacks. For memory efficiency, we employ prototype augmentation which allows us to maintain the decision boundary of previous stages without the need to store any budget explicitly. After each adaptation, we refine the classifier to avoid unnecessary expansion. Simultaneously, a lightweight model is utilized to ensemble the defense model and the target model by estimating reliable logits for input data, ensuring maximum classification accuracy for both clean and adversarial data.
Extensive experiments on CIFAR-10 and ImageNet-100 show that CAD effectively defends against multi-stage attacks with a few-shot defense budget while maintaining high accuracy on clean data. To the best of our knowledge, this is the first exploration of continual adversarial defense and its real-world deployment.
Our main contributions can be summarized as follows:
$\bullet$ To counter evolving adversarial attacks, we put forward a deployment of continual defense strategy with a practical threat model where diverse attacks emerge over stages. The defense system must adapt well to these attacks using a few-shot defense budget and efficient memory, while retaining knowledge of past attacks and ensuring high performance on both clean and adversarial data.
$\bullet$ We propose, for the first time, the Continual Adversarial Defense (CAD) framework that defends against attacks in the dynamic scenario under four practical principles: continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial data. CAD leverage cutting-edge techniques including continual learning, few-shot learning, non-exemplar class incremental learning, and ensemble learning to satisfy the principles.
$\bullet$ Extensive experiments validate the effectiveness of CAD against multiple stages of modern adversarial attacks using few-shot budget and efficient memory and demonstrate significant improvements over baseline methods.
2 Related Work
2.1 Adversarial Defense
Researchers have proposed various general defense methods against adversarial attacks [23, 36, 20, 34]. As a standard defense method, adversarial training [41, 40, 8, 35, 17, 5, 11] aims to enhance the robustness of the target model by training it with adversarial examples. Another branch of adversarial defense involves purifying the data stream [26, 10, 38, 24] to remove potential adversarial perturbations or noise that could deceive the model. To counter attacks in dynamic environments, model adaptation methods [2, 12, 37, 18] continuously adjust their parameters, state, or activations when encountering attacks to enhance robustness.
In fact, all of the above methods have shown limitations in robustness, resulting in poor classification performance on both adversarial and clean data. In this paper, we present the first Continual Adversarial Defense (CAD) framework as a solution to effectively counter a wide spectrum of attacks emerging stage by stage, while ensuring robust performance in classifying clean data.
2.2 Continual Learning
Continual learning (CL) [21, 14, 32, 43] aims to learn from a sequence of new classes without forgetting old ones.In recent years, some methods have been aimed at solving the CL problem without relying on preserving data (none-exemplar) [46], and few-shot scenarios [39] in which only a small number of new class data is gainable. Attempting to break the one-shot defense assumption, a concurrent work [44] is proposed to alleviate catastrophic forgetting in a simple continual defense scenario. In this paper, we convert the proposed defense scenario into a few-shot and non-exemplar CL setting, which necessitates the defense mechanism to utilize few-shot defense budget and efficient memory for adaptation.
3 Threat Model and Defense Principles
The defense system is deployed online and shielding the target model from a dynamic environment marked by a plethora of unique attacks. As time progresses, new attacks accompanied by poisoned examples will emerge on the Internet, aiming to exploit vulnerabilities in the target model and achieve breakthroughs. In consideration of this reality, we propose a practical threat model in which distinct attacks emerge over stages, requiring the defense system to adapt to evolving threats through the utilization of the defense budget (i.e. adversarial examples and their ground-truths). Typically, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget. The defense budget is crucial as it informs and influences the defender’s response to emerging threats, often discovered during screening or cleanup efforts in the real world, and is provided by web crawlers, security services, and users of the target model. The continual defense scenario is formulated as follows.
Figure 2: Illustration of CAD. Treating new attack classes as incremental, we use continual adaptation to protect the target model while preventing catastrophic forgetting. In stage 0, we reserve embedding space during initial training to mitigate future overfitting. A weight estimator model $f_{p}$ is trained for model ensemble. In stage $i$ ($i>0$), we freeze the feature extractor $\varphi$ of $f_{i}$, fine-tune the expanded classifier $g_{c}^{i}$ with a few-shot defense budget, and store class-wise prototypes in a cache for efficient memory use. Finally, $f_{t}$ and $f_{i}$ are ensembled via $f_{p}$ for robust classification.
Our scenario is established on the $N$-way $K$-shot classification paradigm. Attacks are launched under the gray-box setting [28] in which the attacks know the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space. The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at stage 0. At stage $i$, where $i=1,2,\cdots,T$, new attacks $A_{i}(\cdot)$ are detected and the defender receives a set of budget $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ [^1] In the adversarial attack formula $\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x},y,f_{t})$, we omit the target model $f_{t}$ and the ground-truth $y$. containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack, where $y$ is the ground truth of sample $\mathbf{x}$. Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
Taking reality into account, the defense mechanism should satisfy the following principles:
Principle 1
Continual adaptation to new attacks without catastrophic forgetting.
In the dynamic environment, various new attacks will emerge on the internet as time goes on. Therefore, the defense mechanism must be capable of adapting to a range of new attacks across different time stages while also retaining knowledge of previous ones.
Principle 2
Few-shot adaptation.
The increase in defense budget directly correlates with the frequency of successful attacks on the target model. The defense system must proactively adapt the defense based on few-shot budget to prevent potential security disasters that could arise from delayed action.
Principle 3
Memory-efficient adaptation.
Over time, the ongoing influx of attacks results in accumulating defense budgets, potentially leading to memory constraints. In practical terms, the defender may not have sufficient memory space to accommodate this accumulation. Consequently, storing all received defense budget for adaptation is impractical and should be avoided.
Principle 4
High accuracy in classifying both clean and adversarial data.
The previous defense strategies have led to a sacrifice in the classification accuracy of clean data, resulting in performance degradation. This decline could potentially harm crucial real-world business operations. As a result, a defender must prioritize maintaining performance on both clean and adversarial data.
4 Continual Adversarial Defense Framework
As shown in Figure 2, Continual Adversarial Defense (CAD) framework consists of continual adaptation to defend the target model while preventing catastrophic forgetting, embedding reservation to alleviate overfitting brought by the few-shot budget, prototype augmentation for memory efficiency, and model ensemble to ensure the classification performance on both clean and adversarial data.
4.1 Continual Adaptation to New Attacks
In response to Principle 1, we use continual adaptation to defend against new attacks while preventing catastrophic forgetting. At stage 00, an initial defense model $f_{0}:\mathcal{X}\to\mathbb{R}^{N}$ which consists of a feature extractor $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and a classifier $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ is optimized under full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. The defense model is designed to tackle adversarial examples, complementing the target model. At stage $i$, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to new attacks using the defense budget $\mathcal{A}_{\mathrm{train}}^{i}$ while mitigating catastrophic forgetting, as outlined in Principle 1, with cached prototypes $P_{e}$ extracted from past attacks.
Inspired by continual learning (CL) [25], we regard the classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion. Following [21], in stage $i$ $(i>0)$, we expand the classifier $g_{c}^{i-1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $g_{c}^{i}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$. The parameters of the expanded classifier consist of those from the old classifier and newly initialized ones: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$. Meanwhile, the ground truth of $\mathbf{x}_{\mathrm{adv}}^{i}\in\mathcal{A}_{\mathrm{train}}^{i}$ is rewritten corresponding to the incremental classes with the new attack  [^2] We omit the networks’ parameter in formulas.:
$y^{i}=y+N\times{i}.$ (1)
The initial defense model $f_{0}$ will be trained in Section 4.2. After this, we freeze the feature extractor $\varphi$ and fine-tune the expanded classifier part for the new attack $A_{i}$ using the few-shot budget $\mathcal{A}_{\mathrm{train}}^{i}$ in Section 4.3.
Due to the internal similarity of some adversarial attacks [33], trained with one attack, the defense model can, to some extent, generalize to other attacks, resulting in some unnecessary class expansion. Old attacks occurring in new stages also lead to unnecessary adaptation. Therefore, before fine-tuning at each stage, we perform model testing to filter out the useless data from the defense budget that can be correctly classified.
For evaluation, we focus solely on the class assignment of the image, rather than the specific attack it is subjected to. Therefore, the prediction for an instance $\mathbf{x}$ is:
$y_{\mathrm{pred}}=(\arg\max f_{i}(\mathbf{x}))\bmod{N}$ (2)
4.2 Embedding Space Reservation for Few-Shot Adaptation
A problem brought by the few-shot budget in Principle 2 is over-fitting which is common in few-shot learning [27]. To tackle this issue, we choose to reserve embedding space for future attacks by generating and assigning virtual prototypes, squeezing the embedding space of previous attacks.
First, before training the defense model $f_{0}$ at stage 00, several virtual classes $W_{v}=[\mathbf{w}_{v}^{1},\cdots,\mathbf{w}_{v}^{V}]\in\mathbb{R}^{d\times V}$ are pre-assign in the classifier [39], where $V=N\times T$ is the number of virtual classes, i.e. the reserved classes for future attacks. Therefore the output of current model is $f_{0}(\mathbf{x})=[W^{0},W_{v}]^{\top}\varphi(\mathbf{x})$. After training, $W_{v}$ will be used to initiate the new parameters of the expanded classifier.
Second, virtual instances are constructed by manifold mixup [31]:
$\mathbf{z}=h_{2}(\omega h_{1}(\mathbf{x}_{\mathrm{adv},r}^{0})+(1-\omega)h_{1}(\mathbf{x}_{\mathrm{adv},s}^{0})),$ (3)
where $\mathbf{x}_{\mathrm{adv},r}$ and $\mathbf{x}_{\mathrm{adv},s}$ belong to different classes $r$ and $s$, and $\omega\sim\mathrm{B}(\alpha,\beta)$ is a trade-off parameter the same as [43]. $h_{1}$ and $h_{2}$ are decoupled hidden layers of feature extractor i.e. $\varphi(\mathbf{x})=h_{2}(h_{1}(\mathbf{x}))$.
Finally, the embedding space reserving is conducted by training $f_{0}$ with the following loss:
$\mathcal{L}_{0}=$ $\gamma F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0}),\hat{y})+F_{\mathrm{ce}}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0})+$
$+$ $\gamma F_{\mathrm{ce}}(\mathsf{Mask}(f_{0}(\mathbf{z}),\hat{y}),\hat{\hat{y}})+F_{\mathrm{ce}}(f_{0}(\mathbf{z}),\hat{y}).$ (4)
where $\hat{y}=\mathrm{argmax}_{j}\mathbf{w}_{v}^{j\top}\varphi(\mathbf{x}_{\mathrm{adv}}^{0})+N$ is the virtual class with maximum logit, acting as the pseudo label. $\hat{\hat{y}}=\mathrm{argmax}_{k}\mathbf{w}_{k}^{0\top}\mathbf{z}$ is the pseudo label among current known classes. $\gamma$ is a trade-off parameter, $F_{\mathrm{ce}}$ represents the standard cross-entropy loss [42], and function $\mathsf{Mask}(\cdot)$ masks out the logit corresponding to the ground-truth:
$\mathsf{Mask}(f_{0}(\mathbf{x}_{\mathrm{adv}}^{0}),y^{0})=f_{0}(\mathbf{x}_{\mathrm{adv}}^{0})\otimes(\mathbf{1}-\mathrm{onehot}(y^{0}))$ (5)
where $\otimes$ is Hadamard product and $\mathbf{1}$ is an all-ones vector.
In Eq. 4.2, it pushes instance $\mathbf{x}_{\mathrm{adv}}^{0}$ away from the reserved virtual classes (item 1) while towards known classes (item 2), and forces virtual instance $\mathbf{z}$ away from known class (item 3) while towards virtual class (item 4). Trained with $\mathcal{L}_{0}$, the embedding of initial benign classes will be more compact, and the embedding spaces for virtual classes will be reserved [43]. The reserved space allows the defense model to be adapted more easily in the future and alleviates overfitting brought by few-shot budget.
4.3 Prototype Augmentation for Memory Efficiency
In response to Principle 3, we use prototype augmentation to efficiently manage memory.
When learning new classes, the decision boundary for previous classes can be dramatically changed, and the unified classifier is severely biased [45]. Many CL methods store a fraction of old data to jointly train the model with current data to address this issue [25]. However, preserving old budget may also lead to memory shortage. For memory efficiency, we adopt prototype augmentation [46] to maintain the decision boundary of previous stages, without saving any budget. We memorize one prototype in the deep feature space for each class with each attack, and over-sample (i.e., $U_{B}$) prototypes $\mathbf{p}_{B}$ and ground-truths $y_{B}$ to the batch size, achieving the calibration of the classifier:
$\mathbf{p}_{B}=U_{B}(P_{e}),\quad\mathcal{L}_{\mathrm{proto}}=F_{\mathrm{ce}}(\mathbf{p}_{B},y_{B})$ (6)
where $P_{e}=\{\mathbf{p}_{e}^{j}\}_{j=0}^{N\times i}$ is the set of class-wise average embeddings [27], i.e., the prototype cache. After each adaptation, prototypes of the current attack are added into $P_{e}$:
$\mathbf{p}_{e}^{j}=\frac{1}{K}\textstyle{\sum_{k=1}^{|\mathcal{A}_{\mathrm{train}}^{i}|}}\mathbb{I}(y_{k}^{i}=j)\varphi(\mathbf{x}_{\mathrm{adv},k}^{i})$ (7)
where $\mathbb{I}(\cdot)$ is the indicator function, $K$ the amount of budget for each class, and $(\mathbf{x}_{\mathrm{adv},k}^{i},y_{k}^{i})$ the $k$-th sample in $\mathcal{A}_{\mathrm{train}}^{i}$.
At each stage, the defense model is adapted using the following loss:
$\mathcal{L}_{\mathrm{ce}}=F_{\mathrm{ce}}(f_{i}(\mathbf{x}_{\mathrm{adv}}^{i}),y^{i}),\quad\mathcal{L}_{f}=\mathcal{L}_{\mathrm{ce}}+\lambda\mathcal{L}_{\mathrm{proto}}$ (8)
4.4 Model Ensemble for Clean Data Classification
To maintain high classification accuracy on both clean and adversarial data mentioned in Principle 4, we propose the model ensemble as the last part of CAD. Ensemble adversarial training (EAT) [30] which trains a robust model using adversarial samples generated by the target model is a simple yet effective way to defend against adversarial attack under the gray-box setting while maintaining the classification performance of clean data. We extend EAT to our scenario by training a small weight estimator model $f_{p}$ to fuse the logits of models.
We adopt self-perturbation [33] to train the weight estimator at the first stage. Agnostic to any of the attacks, a self-perturbation-trained model is able to distinguish images from benign and adversarial. The weight estimator $f_{p}$ outputs a weight vector $\mathbf{w}\in\mathbb{R}^{2}$ to ensemble the defense model and the target model:
$\mathrm{logit}_{i}(\mathbf{x})=\mathbf{w}\cdot[f_{t}(\mathbf{x}),f_{i}(\mathbf{x})]^{\top}$ (9)
In this way, both clean and adversarial images are correctly classified.
For the overall algorithm of the CAD framework, details about self-perturbation, and limitations of CAD, please refer to the supplementary materials.
5 Experiment
To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies comparing CAD with baselines from three research streams. For evaluation and analysis, we use two metrics in this section: (1) Classification accuracy (Acc) after each adaptation for defense performance evaluation. (2) Average accuracy of all attacks occurring within and before each stage, i.e. average incremental accuracy (AIAcc) [25], to assess the model’s knowledge retention across all attacks within and before each stage.
Method Classification accuracy(%) against $l_{\infty}$ attacks at each stage. Clean
0:PGD 1:SNIM 2:BIM 3:RFGSM 4:MIM 5:DIM 6:NIM 7:VNIM 8:VMIM
None-defense 0.2 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.0 89.8
TRADES [40] 46.5 45.5 49.6 46.2 48.0 47.7 46.7 48.2 49.5 76.2
CLC [2] 48.4 48.2 49.8 47.3 49.6 48.8 47.9 49.1 50.5 77.5
GAIRAT [41] 59.8 60.2 58.4 59.2 58.8 58.9 58.9 59.7 59.8 74.1
DMAT [35] 64.3 65.1 66.2 65.9 65.4 64.4 64.1 65.8 65.6 75.9
MeanSparse [1] 69.4 69.3 70.5 70.1 71.3 71.8 70.0 71.6 71.4 78.8
ADP [38] 77.2 76.9 71.8 71.8 75.6 75.9 76.2 76.0 72.3 74.1
DiffPure [24] 53.2 53.0 52.4 55.8 53.9 53.4 52.6 54.7 52.3 73.6
CAD(ours) 83.7 83.4 82.0 82.5 83.5 82.0 81.9 82.2 81.9 88.1
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 83.5 83.9 84.0 82.6 84.3 83.8 84.2 82.2 81.4 87.8
Table 1: Adversarial defense performance on ImageNet-100. Volumn ”Clean” denotes standard accuracy on clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ represents premium-CAD without memory or defense budget constraints, serving as the empirical upper bound (gray background).
Method Average incremental accuracy(%) against $l_{p=0,1,2,\infty}$ attacks at each stage.
0:PGD 1:SNIM 2:VNIM 3:CW 4:DeepF. 5:EAD 6:EADEN 7:OnePixel 8:SparseF.
None-defense 0.6 1.3 0.0 0.0 8.6 0.0 0.2 39.1 29.7
EWC [14] 95.5 68.2 67.0 23.7 23.6 23.7 26.9 25.9 22.2
SSRE [46] 89.1 82.5 78.1 65.8 58.4 50.6 47.5 42.4 37.8
BEEF [32] 94.9 83.1 84.7 82.3 77.9 74.2 73.8 72.6 64.5
FACT [43] 94.5 87.2 86.4 81.5 80.9 78.6 78.4 78.2 74.6
CAD(ours) 94.5 89.1 90.5 84.1 84.8 81.1 82.1 82.8 78.2
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.1 95.2 94.9 95.3 94.6 96.0 95.2
Table 2: Continual defense performance on CIFAR-10. DeepF. and SparseF. denote DeepFool and SparseFool respectively.
Datasets
Experiments are conducted on CIFAR-10 and ImageNet-100. CIFAR-10 [15] contains 5000 images for training and 1000 images for testing in each of the 10 classes. ImageNet-100 [29] is a subset of ImageNet [4] and contains 100 classes with 1000 images for training and 100 images for testing in each class. Images from CIFAR-10 and ImageNet-100 are resized to $32\times 32$ and $224\times 224$ respectively. Before training, the above datasets endured data augmentation including horizontal flip and random crop.
Attacks
We choose PGD-$l_{\infty}$ [23] as the initial attack. After that, we select 8 adversarial attack methods under $l_{\infty}$: SNIM [20], BIM [16], RFGSM [30], MIM [6], DIM [36], NIM [20], VNIM and VMIM [34] to compose the attack pool in defense performance evaluation. The perturbation magnitude is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following [40]. To evaluate CAD’s effectiveness in mitigating catastrophic forgetting, $l_{\infty}$ attacks (VMIM and SNIM), $l_{2}$ attacks (CW and DeepFool), $l_{1}$ attacks (EAD and EADEN), and $l_{0}$ attacks (OnePixel and SparseFool) compose the attack pool. Details of attack parameters are given in the supplementary materials.
Implementation Details
Both the target and defense models use ResNet [9] as the backbone: WideResNet-28-10 for CIFAR-10 following [8] and ResNet-50 for ImageNet-100 following [17]. A 4-layer ConvNet is used as the weight estimator model. The cosine classifier [7] is adopted as the classifier of the defense model. We utilize Torchattacks [13] to generate adversarial images. The defense budget per class is set to $K=10$, with $T=8$ stages since there are 9 attacks (PGD as the initial attack). Following [43], Beta distribution parameters are $\alpha=\beta=2$, and trade-off parameters are $\gamma=0.01$. Scaling hyperparameter of prototype augmentation is set to $\lambda=0.5$. Epochs for training and fine-tuning are set to $100$ and $4$ respectively.
Defense Baselines
We compare CAD with robust models TRADES [40], CLC [2], GAIRAT [41], DMAT [35], and MeanSparse [1], and purification methods ADP [38], DiffPure [24] in defense performance. Robust models are evaluated in a black-box setting, while purification methods and CAD are tested in a gray-box setting. To assess CAD’s ability to mitigate catastrophic forgetting, we compare it with CL methods EWC [14], SSRE [46], BEEF [32], and FACT [43] in a gray-box setting. Additionally, we propose a variant of CAD, premium-CAD, as the empirical upper bound, which has no memory or defense budget constraints. In premium-CAD, attack-specific defense models are trained with abundant defense budget and ensembled with the target model via a weight estimator. Details of the baseline and premium-CAD implementation can be found in the supplementary materials.
(a) PGD
(b) CW
(c) EAD
(d) OnePixel
Figure 3: T-SNE diagrams of features of test data encoded by the defense model after all stages. Each color represents a class of CIFAR10. After several steps of adaptation, CAD retains the capability to categorize the attacked data.
5.1 Comparisons to Baselines
To evaluate the defense performance, we first compared CAD to defense baseline methods in defending against 9 adversarial attacks. For CAD, the defense environment is dynamic, with attacks occurring at different stages. The other baseline methods operate in a static defense environment that does not differentiate the order of attacks. As depicted in Table 1, despite the small scale of the defense budget, CAD demonstrates strong adaptability to each type of attack. When compared with the top baseline MeanSparse [1] and ADP [38] on ImageNet-100, CAD outperforms them by approximately $6\%$ for each attack. This underscores its capacity for continuous adaptation in response to emerging attacks. Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $88.1\%$ accuracy, while the target model stands at $89.8\%$. This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images.
Ablation 0:PGD 1:CW 2:EAD 3:OnePixel Clean
w/o $\mathcal{L}_{0}$ 94.5 82.4 76.1 71.8 95.8
w/o $\mathcal{L}_{\mathrm{proto}}$ 94.4 89.5 80.0 78.5 95.8
w/o $f_{p}$ 94.5 89.9 86.6 85.3 55.9
CAD 94.5 89.9 86.6 85.3 95.8
Table 3: AIAcc(%) on CIFAR-10 for ablation.
To assess the effectiveness of CAD in retaining knowledge, we compare it to CL baseline methods on CIFAR-10 in a dynamic scenario formed by attacks under different norms $l_{p=0,1,2,\infty}$. As depicted in Table 2, CAD maintains robust defense performance despite changes in attack types. However, as the variety of attack types increases, the defense performance gradually diminishes. This indicates that defending against an expanding array of attack types presents a more challenging task for continual defense. Compared with the top baseline method FACT, CAD consistently outperforms it by an average of approximately $3.4\%$, highlighting CAD’s effectiveness in mitigating catastrophic forgetting.
It is noteworthy that premium-CAD surpasses all baseline methods. Given sufficient defense budget and memory resources, premium-CAD emerges as the superior choice.
5.2 Ablation Study
As argued, embedding reservation facilitates smoother adaptation and reduces overfitting, prototype augmentation preserves decision boundaries from previous stages, and model ensembling ensures robust classification for both clean and adversarial images. To validate the contribution of each component, we conduct an ablation study using four attacks: PGD ($l_{\infty}$), CW ($l_{2}$), EAD ($l_{1}$), and OnePixel ($l_{0}$). As shown in Table 3, removing $\mathcal{L}_{0}$ and $\mathcal{L}_{\mathrm{proto}}$ results in performance drops of $13.5\%$ and $6.8\%$, respectively, in the final stage. Additionally, removing model ensembling (w/o $f_{p}$) severely degrades performance on clean images. These results confirm the effectiveness of each component.
Figure 4: Catastrophic forgetting verification of CAD and the vanilla model under the attack sequence in Table 2 on CIFAR-10.
5.3 Adaptation Without Forgetting
As stated in Principle 1, a defense model must adapt to new attacks while preventing catastrophic forgetting. To evaluate CAD’s adaptability, we present T-SNE visualizations of features extracted from adversarial test data in Figure 3. These features are encoded by the defense model after the final stage of the attack sequence in Table 2. From (a) to (d) in Figure 3, the feature clusters become increasingly indistinct, indicating that the difficulty of few-shot adaptation grows as new attack types are introduced. Despite this, CAD successfully retains its defense performance against previous attacks, demonstrating its ability to adapt while mitigating catastrophic forgetting.
Furthermore, we compare CAD with a vanilla model as the empirical lower bound. As shown in Figure 4, without CAD, the model’s forgetting of previous attacks worsens over stages,with its error rate on PGD and CW exceeding 50% by stage 8. This highlights CAD’s effectiveness in preventing catastrophic forgetting.
5.4 Number of Shots for Adaptation
$K$ 0:PGD 1:CW 2:EAD 3:OnePixel
$15$ 94.5 90.0 88.4 85.5
$10$ 94.5 89.9 86.6 85.3
$5$ 94.5 89.4 86.1 84.6
$1$ 94.5 87.9 84.6 83.4
Table 4: AIAcc(%) at different budget scale $K$.
As indicated in Principle 2, the defense adaptation relies on few-shot defense budget. To investigate the influence of the budget shots $K$, we conducted tests with $K=15,10,5,1$ on CIFAR-10. The results presented in Table 4 illustrate that defense performance diminishes as $K$ decreases. Nevertheless, even with reduced $K$, CAD continues to efficiently defend against adversarial attacks ($83.4\%$ when $K=1$ at stage $3$), highlighting its strong capability in few-shot scenarios.
Method Cache Size per Stage Model Size per Stage
1 2 3 1 2 3
CAD 50.3K 75.5K 100.4K 141M 141M 141M
$\mathrm{\textbf{CAD}}^{\dagger}$ 257M 377M 504M 280M 419M 558M
Table 5: Comparison of the Memory cost (Byte) between CAD and premium-CAD.
5.5 Memory Allowance for Adaptation
As outlined in Principle 3, efficient memory usage is crucial for continual defense. To evaluate CAD’s efficiency, we measure memory footprint and model size at each stage on CIFAR-10. As shown in Table 5, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size. This underscores our approach’s ability to utilize efficient memory.
5.6 Defense Against White-Box Attack Sequence
To further evaluate the robustness of CAD, we conducted white-box experiments on CIFAR-10. The white-box setting follows the same configuration as AIR [44], where attacks occur in real-time during both adaptation and testing, and there is no restriction on the amount of defense budget used for adaptation. Moreover, both the target model and the defense model are fully visible to the attacker. To adapt to white-box attacks, the feature extractor $\varphi$ is no longer fixed but learnable.
We compare CAD with vanilla adversarial training (AT) and CAD + robust model under a white-box attack sequence. Here, CAD + robust model refers to using a robust model [35] as the backbone in CAD. As shown in Table 6, CAD outperforms vanilla AT in adapting to new attacks and further strengthens the robust model’s defense against evolving threats.
Method 0:PGD 1:SNIM 2:BIM 3:RFGSM
Vanilla AT [22] 36.2 35.8 34.5 34.2
CAD 51.7 55.2 53.6 52.5
$\mathrm{\textbf{CAD}}^{\dagger}$ 51.7 55.4 53.8 53.4
Robust Model [35] 72.1 78.8 71.7 71.6
CAD + Robust Model 74.5 79.6 73.4 73.8
$\mathrm{\textbf{CAD}}^{\dagger}$ + Robust Model 74.5 81.2 78.8 78.9
Table 6: Classification accuracy (%) against white-box attacks.
5.7 Comparison with AIR
We compare CAD’s performance in the white-box setting with the results reported in the AIR paper. As shown in Table 7, after adapting from Attack 1 to Attack 2, CAD better retains its defense against previous attacks compared to AIR. This is attributed to CAD’s prototype augmentation mechanism, albeit at the cost of a slightly larger prototype cache. Additionally, since premium-CAD trains a separate defense model for each attack, it completely avoids forgetting previous attacks after adaptation.
Method FGSM to PGD PGD to FGSM
Attack 1 Attack 2 Attack 1 Attack 2
AIR [44] ${53.4}^{\ddagger}$ ${44.1}^{\ddagger}$ ${43.0}^{\ddagger}$ ${52.3}^{\ddagger}$
CAD 54.2 48.6 47.5 53.9
$\mathrm{\textbf{CAD}}^{\dagger}$ 55.8 51.7 51.7 55.8
[…]
6 Conclusion
The operational environment for defense systems is inherently dynamic, making it unrealistic to expect any defense method to effectively address all types of attacks. In response to this challenge, we introduce the first Continual Adversarial Defense (CAD) framework. CAD is engineered to collect defense budget from the Internet and dynamically adapt to a diverse range of attacks as they emerge in sequential stages. In consideration of practical applicability, we have formulated four principles that include continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial data, and CAD demonstrates strong performance when operating under these principles. Through extensive experiments, we have demonstrated the effectiveness of our approach in combating multiple stages of various adversarial attacks, achieving significant improvements over baseline methods.